Solve the problem below.


The average monthly salary for a newly hired teacher in a private school is P18,000. If the hiring salary at this school is normally distributed with a standard deviation of P3,000, what is the probability that a newly hired employee randomly selected from a list receives a monthly salary which is less than P15,000?


Write your solution here:


1
Expert's answer
2021-03-26T10:29:42-0400

We have that:

"\mu=18000"

"\sigma=3000"

"P(X<15000)=P(Z<\frac{x-\mu}{\sigma})=P(Z<\frac{15000-18000}{3000})=P(Z<-1)=0.1587"


Answer: the probability that a newly hired employee receives a monthly salary less than P15,000 is 15.87%.